OpenStudy (anonymous):

Multiply and simplify by factoring: forth-root(16)*forth-root(128)

OpenStudy (anonymous):

\[\sqrt[4]{16}\times \sqrt[4]{128}\] is how it would really look

OpenStudy (anonymous):

ok \[16=2^4\] \[128=2^7\]so you have \[\sqrt[4]{2^{11}}\] 4 goes in to 11 twice leaving a remainder of 3, giving \[2^2\sqrt[4]{2^3}=4\sqrt[4]{8}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw